OpenStudy (anonymous):

Write in vertex form y=x²-9x pleasee?

OpenStudy (saifoo.khan):

\[(x-\frac92)^2-\frac{81}{4}\]

OpenStudy (anonymous):

h= 4.5 or 9/2 k= 20.25 or 81/4 Vertex form = (x-9/2)^2 - 81/4